If X is the number scored in a throw of a fair die, show that the Chebychev’s
inequality gives P{ |X −µ| >2.5 } < 47.0 , where µ is the mean of X, while the actual
probability is zero.
1
Expert's answer
2020-03-10T12:20:11-0400

Chebyshev's inequality:

"P\{|X-\mu|\geq k\sigma\}\leq\frac{1}{k^2}, k\in R."

Let X is our random variable.

"MX=\frac{1}{6}(1+2+3+4+5+6)=3.5.\\
MX^2=\frac{1}{6}(1^2+2^2+3^2+4^2+5^2+6^2)=\frac{91}{6}.\\
DX=MX^2-(MX)^2=\frac{91}{6}-(3.5)^2.\\
\sigma X=\sqrt{\frac{91}{6}-(3.5)^2}.\\
k\sigma X=2.5.\\
k=\frac{2.5}{\sqrt{\frac{91}{6}-(3.5)^2}}\approx 1.46385.\\
\frac{1}{k^2}\approx 0.4666."